OpenStudy (anonymous):

how do you Find the slope of a line that passes through (–2, –3), and (1, 1)

OpenStudy (anonymous):

slope=(y1-y2)/(x1-x2) =(1-(-3))/(1-(-2)) =(1+3)/(1+2) =4/3

OpenStudy (turingtest):

Please post this in the math section. Thanks.

OpenStudy (anonymous):

@msumner

OpenStudy (anonymous):

thanks but my teacher helped me out